Investors require a 15% rate of return on Brooks Sisters' stock (rs = 15%).

  1. What would the value of Brooks's stock be if the previous dividend was D0 = $1 and if investors expect dividends to grow at a constant compound annual rate of (1) - 2%, (2) 0%, (3) 5%, or (4) 13%? Round your answers to the nearest cent.
    1. $    
    2. $    
    3. $    
    4. $   
Answer #1

Current price=D1/(Required return-Growth rate)

1.Current price=(1*(1-0.02)/(0.15-(0.02)

=0.98/0.17=$5.76(Approx)

2.Current price=1/0.15

=$6.67(Approx)

3.Current price=(1+0.05))/(0.15-0.05)

=$10.5

4.Current price=(1+0.13)/(0.15-0.13)

=$56.5